OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x2 - 3 and g(x) = √(3+x)

OpenStudy (anonymous):

please help!!! 15 mins to complete the quiz!! 2nd to last question left!!

OpenStudy (anonymous):

substitute g(x) for x in f(x) f(x) = x2 - 3 f(g(x)) = ( sqrt 3+x)^2 - 3 f(g(x)) = 3+x-3 solve this u get ur answer similar way u do it for g(f(x))

OpenStudy (anonymous):

hmm.. so itd be = x+6? im terrible at math and dont understand this question at all

OpenStudy (anonymous):

noo ... first to find f(g(X) substitute g(x) in place of x tell me what to do get???

OpenStudy (anonymous):

hmmm..

OpenStudy (anonymous):

idk i get comfused with the square root sign

OpenStudy (anonymous):

ok   got it ???